HUBBLE SPACE TELESCOPE DAILY REPORT #922 PERIOD COVERED: 0800 EST Daily Status Report as of 184/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.2 Completed PC 4746 (Investigation of the Dynamics of Binary Asteroids) The Planetary Camera was used to make four images of the asteroid KLOPA-216 which is suspected to be a binary asteroid based on its unusual light curve. All four observations were received and displayed. They were reported to look okay. 1.3 Uplinked Updated Ephemeris Tables, Per 2871 1.4 Completed PC 4714 (WF/PC Cycle 2 Calibration: Internal Delta Flats) The Faint Object Camera was used to make four internal flat calibration images which will be used for the restoration of Earth flat images, which were altered due to changes in quantum efficiency which occurred after decontamination operations. The observations were received and displayed in the OSS. They were reported to look okay. 1.5 Completed Two Sets of WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. A total of three image observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.6 Completed GHRS/2 3759 (Dust, Molecular, Atomic Composition and the Physical Condition in Mol.) The High Resolution Spectrograph was used to make three spectral observations of standard star HD102965 which lies behind a molecular cloud. The absorption lines in the spectra of this star will provide information on the gas and dust in the foreground cloud. All three observations were received and displayed in the OSS. No anomalies were noted. 1.7 Completed GHRS/2 4580 (High Resolution Observations of the Jovian UV Aurora - Cycle 3 Med.) The High Resolution Spectrograph was used to make two kinetic temperature measurements of the Jovian auroral regions, to better understand Jupiter's upper atmosphere. The observations were of the ultraviolet Lyman and Werner bands in the 1586 to 1620 A0 region. An onboard acquisition was performed on the Moon Io , and then the telescope was offset to Jupiter's auroral region. The observations were executed as scheduled, and no problems were reported. Both spectral observations were reported to look okay. 1.8 Completed GHRS/2 3614 (Boron as a Probe of Stellar Structure and Mass Loss) The High Resolution Spectrograph was used to make 12 spectral observations of the Boron lines in stars HD128620 and HD128621. The abundance of Boron can be used to study the process on the interior of these stars. All spectral observations were reported to look okay. 1.9 Completed PC 3230 (PC PSF and Focus Monitoring 2) The Planetary Camera was used to make one set of images which will be used to monitor the point spread function and focus changes which have occurred due to changes in desorption and other changes in the WFPC system. The observations were executed with no apparent problems. The star appeared near the fiducial point of each chip. 1.10 Completed GHRS/2 4650 (A Study of Fast Winds in Central Star of Planetary Nebulae: Cycle 3 Med.) The High Resolution Spectrograph was used to make observations of the central stars of planetary nebulae ABELL78 and K1-16 (Both located in Centaurus) to learn more about the velocities in the winds produced by this stars and the rate of their mass loss. Five observations were received and displayed. The wavelength and spectral observations were reported to look okay. 1.11 Completed WFC 4713 (WF/PC Cycle 2 Calibration: WFC Internal Delta Flats) The Wide Field Camera was used to make a total of five internal flat calibrations images. All five images were reported to be normal. 1.12 Completed Two Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make a snapshot image of the protoplanetary system V354-LAC to study its current morphology. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.13 Completed WFC 3931 (Imaging of Distant Active Galaxies Cycle 2) The Wide Field Camera was used to make two images of the low redshift quasar 1302-102 which had been partially resolved from ground observations. Both observations were executed in fine lock using FGS #1 only. No recentering was reported to have occurred. The observations were received and displayed, and no anomalies were reported. 1.14 Completed WFC 3857 (Morphology of Galaxies in Clusters at Z = 0.5 - Continuation of 2373) The Wide Field Camera was used to make two images to study the morphology of galactic cluster NGC-188 using filter F702W. The observations were executed as scheduled, and no problems were reported. Both images were reported to look okay. 1.15 Completed Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies - FOC/96) The Faint Object Camera was used to complete one snapshot ultraviolet observation of the nearby galaxy ESO-0411-5. The exposures lasted for ten minutes and used filter F220W, and were executed in fine lock. No anomalies were noted. 1.16 Completed FOC/96 3801 (Search for QSOs suitable for Subsequent Observation of He II) The Faint Object Camera was used to make observations of quasar 2318+0119 to measure the levels of its ultraviolet flux, to see if the flux levels are high enough for subsequent observations using the GHRS and FOS. The observations were executed in fine lock during LOS. The observations were later received and displayed in the OSS. No anomalies were noted. 1.17 Completed PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor - Part 2) The WFPC was used to make four internal flat calibration images to check for new forms of contamination (i.e., Daises and measles), and to check for the possible return of quantum efficiency hysteresis. The observations completed as scheduled, and no problems were reported. 1.18 Completed GHRS/2 4476 (Deuterium Observations of Jupiter) The High Resolution Spectrograph was used to make six medium and high resolution observations profiles of the D-Lyman line emitted at the limb of the Jovian atmosphere. All six observations were reported to look okay. The target flux levels were reported to be 1171. 1.19 Complete 4113 (Auroral Observations of Jupiter) The Faint Object Camera was used to make two images of the auroral emissions on the planet Jupiter at 1600 A0. Both north and south observations were contained within the sequence. The observations were received, and reported to look okay. 1.20 Completed WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of four observations were received. One observation was reported to be saturated, and the others were okay. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 25 Successful : 25 Scheduled Reacquisitions : 29 Successful : 29 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 44 Successful : 43 184/2205Z the FHST 2 and 3 failed to perform a full maneuver update (secondary). This did not cause a Guide Star Acquisition Failure. 2.3 Significant HSTARs: None 2.4 Operations Note: The MSS select flag (QDSMSSU) toggled from a value of 2 to a value of 1 on Day 186/0129Z for approximately three minutes. The out of limit period occurred while the vehicle was executing a type two slew. HSTAR #4203 was written because the out of limit did not occur during a SAA event (weak magnetic field). PCS is currently analyzing sensor temperature, magnetic field strength data, and the MSS resonableness test logic. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flat Narrow. 2) 3.2 Complete WFC 4381 (HST Medium Deep Survey - Cycle 3) 3.3 Complete FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: Cycle 3 High) 3.4 Complete FGS 5055 (A critical Test of the Galactic Escape Velocity at R(Sun): Cycle 1 Obsv.) 3.5 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.6 Complete FOC/96 4969 (Search for QSOs Suitable for Subsequent Observation of He II) 3.7 Complete GHRS/2 4476 (Abundance and Distribution of Atomic Deuterium in the Jovian Upper Atmosphere) 3.8 Complete PC 4929 (WF/PC Cycle 2 : Improved Dark Calibration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-